User name Roedelbronn
Log entry time 08:12:44 on February17,2002
Entry number 82933
keyword=Owl shift summary
00:17 Started run 2557 (p = 2.619)
00:31 Started run 2558 (p = 2.580) The last run was only approximately 6 minutes long, but it took a long time for the HRS dipole to ramp to the new current.
00:46 Started run 2559 (p=2.541)
00:59 Started run 2560 (p=2.503)
01:11 Started run 2561 (p=2.464)
01:25 Started run 2562 (p=2.425)
01:32 Shut down beam in preparation for Controlled Access to flip polarity of HRS dipole.
02:40 Ed and I finished reversing the polarity of the HRS dipole. We switched to the optics target.
03:39 Took a long time to get the HRS dipole magnet to desired field. Starting 3A(b) optics run
04:00 Beam down for daily quantum efficiency measurement
04:30 Switched to LH2 norad target
07:25 Beam is down for an accelerator IOC reboot